Database last updated on Wed, 28 Jan 2026 20:40:51 +0000 by localhost
Instrument constant changes for the week before 2025050 (2 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 2 rows)
HG burned out? (AP results when HG is ON) (0 sec, 1 rows)
Non-critical warnings from b-files' comments for days between 2025050 and 2025050 (0 sec, 1 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 050 05 44 51 Reset Warm -> Tepid from 2025 050 05 44 45 until 2025 050 05 44 45 . Counts: 1,0 | 1 | 3 | B05025.192 | TORON | #192 | 05:37:31 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | aode | ds | dt | ed | fm | fz | hg | hp | rs | sl | uv | zs |
| 050 | TORONTO | #192 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.21) | 47 | 47 | 2 | 2 | 19 | 9 | 12 | 12 | 2 | 7 | 23 | 36 |
(0 sec)
Relative and absolute humidity for day 2025050 (2 sec, 1 rows)
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
| 192 | 8.4 | 1.3 |